OpenStudy (anonymous):

x2 - 45x + 324 = 0 middle term splitting method?

OpenStudy (ajprincess):

Can u find two numbers a and b such that a+b=-45 a*b=324?

OpenStudy (unklerhaukus):

\[324=2\times162=2^2\times81=2^2\times3^4\]

OpenStudy (anonymous):

also can you tell me the technique to split high value integers

OpenStudy (anonymous):

@UnkleRhaukus it is only factorization of 324. how will it help to split the terms??

OpenStudy (anonymous):

@ajprincess i can find the p(x) with product and sum of roots

OpenStudy (unklerhaukus):

\[2\times2\times3\times3=36\] \[3\times3=9\]

OpenStudy (anonymous):

thank you i understand now..:)

OpenStudy (unklerhaukus):

\[x^2 - 45x + 324 = 0 \] \[\left(x^2-(36+9)x+(36\times9)\right)=0\]

OpenStudy (anonymous):

thank you everybody for helping me and also for taking a look too :) bye!!

Directrix (directrix):

@Ekaansh The technique is known as "busting the b," the name coming from the b coefficient of the linear x term in the expression ax^2 +bx + c where a is not 0. Follow this link to several examples of busting the b: http://mysite.cherokee.k12.ga.us/personal/randy_smith/site/Lists/Math%203%20Class%20Notes/Attachments/78/Alg%20Rev%20C%20Day%203%20notes.pdf

OpenStudy (unklerhaukus):

if the numbers are nice enough and small enough , this technique can be done in ones head., In all other cases i would just use the quadratic formula because i can always remember it and it always works , (even for complex solutions) , But If you trying the 'busting the b' method, it is a good idea to factorise the large numbers